OpenStudy (anonymous):

What is 12.5 of what number is 35

OpenStudy (imstuck):

12.5 x = 35. 12.5 of what number is translated to 12.5 times x and the word "is" means equals. So 12.5x=35 x=2.8

OpenStudy (campbell_st):

well 12.5% of x = 35 so change it to a fraction \[\frac{12.5 \times x}{100} = 35\] make x the subject... then \[12.5 \times x = 3500\] \[x = 3500 \div 12.5\] hope it helps an alternate method is change the percentage to a decimal 0.125x = 35 x = 35/0.125 bith give the same answer